# Generative AI and Firm Productivity:

### v2

@@ Appendix F Details on the Heterogeneous Treatment Effects Across Products
Table F1 reports the product-level heterogeneous treatment effects of the Pre-sale Service Chatbot experiment. We detect tail products (Column 4) benefiting more from GenAI adoption. The results for market concentration (Columns 1 and 2) and price levels (Columns 5 and 6) tend to be nuanced.
@@
Observation 44,614 44,614 44,614 44,614 44,614 44,614
@@
    We classify products into high and low groups based on three key pre-experiment dimensions. (1) Category market concentration, measured by the sales share of the top 1% of products (ranked by annual sales) within each category. Products in the low group belong to categories with concentration levels below the platform average. (2) Annual sales quantity, also defined within the category. Products in the low group—referred to as tail products—are those comprising the bottom 50% cumulative share of total units sold when ranking products by annual sales quantity within each category. (3) Product price, defined within each category to control for category-level pricing variation. Low-priced products are those priced below the median of their respective category.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
Search Query Refinement.
In Table F2, we find that the gains are only significant for products in low-concentration categories (Column 2), products with fewer sales volume (Column 4), and products with high price level (Column 5).
@@
Observations 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382
@@
    Please refer to Table F1 for detailed notes.
Product Description.
In Table F3, for market concentration and price level, we only observe significant gains for high-concentration categories (Column 1) and high-priced products (Column 5). The conversion rate increase is only significant for head products, while the sales increase is significant for both head and tail products (Columns 3-4).
@@
Observations 4,772,937 4,772,937 4,772,937 4,772,937 4,772,937 4,772,937
@@
    Please refer to Table F1 for detailed notes.
Marketing Push Message.
Table F4 presents that tail products (Column 4) and high-priced products (Column 5) receive a significant increase in conversion rate (Panel B). In terms of the market concentration, products in low-concentration categories experience a significant increase in conversion rate from treated consumers (Column 2).
@@
% Change 2.9% 3.0% 0.8% 7.0% 4.1% 2.7%
Observations 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528
@@
    Please refer to Table F1 for detailed notes.
Google Advertising Titles.
In Table F5, Panels A and B report the heterogeneity in Sales and conversion rate, respectively, while all results are insignificant.
@@
% Change -4.1% -2.1% -2.9% -4.4% -4.2% -2.8%
Observations 714,642 529,374 714,642 529,374 714,872 529,144
@@
    Please refer to Table F1 for detailed notes.

### v3

@@ Appendix F Details on Product Treatment Effect Heterogeneity
Table F1 reports the product-level heterogeneous treatment effects of the Pre-sale Service Chatbot experiment. When focusing on conversion rates, we detect statistically significant differences between product groups classified by annual sales quantity, indicating that the GenAI-driven chatbot service is more effective at converting consumers for tail products.
@@
P-value 0.3444 0.0182 0.4139
Panel C: Cart Value
Treat -1.3 -1.108 -0.366 -1.990 -0.258 -1.584*
(1.551) (1.182) (1.370) (1.472) (2.182) (0.926)
% Change -4.22% -4.47% -1.38% -6.86% -0.725% 6.84%
P-value 0.9704 0.4498 0.3981
Observation (Panel A&B) 44,614 44,614 44,614 44,614 44,614 44,614
Observation (Panel C) 1,802 1,274 1,223 1,853 1,151 1,925
@@
    We classify products into high and low groups based on three key pre-experiment dimensions. (1) Category market concentration, measured by the sales share of the top 1% of products (ranked by annual sales) within each category. Products in the low group belong to categories with concentration levels below the platform average. (2) Annual sales quantity, also defined within the category. Products in the low group—referred to as tail products—are those comprising the bottom 50% cumulative share of total sales quantity within each category. (3) Product price, defined within each category to control for category-level pricing variation. Low-priced products are those priced below the median of their respective category.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise. “Cart Value” refers to the expenditure per consumer, conditional on making a purchase.
@@
    Standard errors are in parentheses. % Change is calculated by dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
Search Query Refinement
In Table F2, we find that for sales, the gains are larger and statistically significant only for products in low-concentration categories (Column 2), products with fewer sales volume (Column 4), and products with high price level (Column 5). The differences between these two products groups are statistically significant when products are classified by market concentration and annual quantity, indicating that GenAI reduces search frictions in settings with greater product differentiation or limited sales history.
@@
P-value 0.0888 0.0064 0.7493
Panel C: Cart Value
Treat -0.292 1.106** -0.098 0.587 0.717 -0.032
(0.291) (0.534) (0.319) (0.411) (0.615) (0.132)
%Change -1.49% 4.35% -0.55% 2.46% 2.04% -0.28%
P-value 0.0208 0.2052 0.2583
Observations (Panel A&B) 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382 1,849,382
Observation (Panel C) 101,975 82,912 84,799 109,265 82,272 108,485
@@
    Refer to Table F1 for detailed notes.
Product Description
Table F3 shows that, for sales, statistically significant differences across product groups arise only when products are classified by price level. This pattern is consistent with GenAI being more effective at reducing information asymmetries in settings with higher decision stakes.
@@
P-value 0.5014 0.0845 0.0047
Panel C: Cart Value
Treat 0.186** -0.019 0.018 0.138 0.092 0.032
(0.089) (0.141) (0.108) (0.104) (0.204) (0.071)
%Change 1.72% -0.16% 0.16% 1.24% 0.59% 0.33%
P-value 0.1895 0.4258 0.8605
Observations (Panel A&B) 4,772,937 4,772,937 4,772,937 4,772,937 4,772,937 4,772,937
Observation (Panel C) 133,779 83,687 106,252 115,810 57,458 159,492
@@
    Refer to Table F1 for detailed notes.
Marketing Push Message
Table F4 shows that, in most cases, we do not observe statistically significant differences between the two product groups.
@@
% Change 1.98% 7.60% 2.90% 2.97% 4.06% 2.68%
P-value 0.8299 0.1826 0.1485
Panel C: Cart Value
Treat -0.108 -0.730 -0.566 -0.0896 -2.292* 0.285
(0.504) (1.038) (0.772) (0.546) (1.381) (0.444)
% Change -0.71% -4.24% -3.71% -0.57% 8.92% 2.17%
P-value 0.6095 0.6106 0.0825
Observations (Panel A&B) 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528 13,715,528
Observations (Panel C) 18,432 3,993 5,498 16,927 4,483 17,942
@@
    Refer to Table F1 for detailed notes.
Google Advertising Titles
Table F5 shows no statistically significant differences between the two product groups.
@@
% Change -4.1% -2.1% -2.9% -4.3% -4.1% -2.8%
P-value 0.7661 0.8272 0.8407
Panel C: Cart Value
Treat -0.819 -0.723 -0.389 -1.477 -1.546 0.0417
(1.258) (1.608) (1.221) (1.700) (1.567) (1.159)
% Change -2.48% -2.07% -1.18% -4.18% -3.69% 0.16%
P-value 0.9439 0.6157 0.4577
Observations (Panel A&B) 714,642 529,374 714,642 529,374 714,872 529,144
Observations (Panel C) 2,906 1,905 3,132 1,679 2,347 2,464
@@
    Refer to Table F1 for detailed notes.